OpenStudy (anonymous):

The yearbook committee is distributing yearbooks to students by class. Each class has approximately the same number of students. The table shows the number of volunteers who helped to distribute the yearbooks to each of the four classes and the time needed to distribute the yearbooks. Which function models the data? Number of volunteers (n) 2 5 8 11 Time in minutes (t) 220 88 55 40 nt = 80 nt = 180 nt = 222 nt = 440

OpenStudy (anonymous):

whats nt is that n*t ?

OpenStudy (anonymous):

I believe it is n*t, That is exactly how it is written in the answers, I didn't forget to input a symbol or anything like that.

OpenStudy (anonymous):

@jim_thompson5910 @Luis_Rivera @Mertsj @phi @robtobey @UnkleRhaukus Someone please! I have been bumping this for too long now.

OpenStudy (kropot72):

For each class, multiply the number of voluteers by the number of hours to distribute the books. You will find the volunteer-hours (nt) to be the same for the distribution to each class. This gives you the ability to make the correct choice of answer.

OpenStudy (kropot72):

@kirk.freedman Do you follow?

OpenStudy (phi):

number of volunteers who helped to distribute the yearbooks to each of the four classes This is a "think about it" question. If you have 2 people handing out n books takes 220 minutes. How much time would it take for just 1 person? I would say twice as long or 440 minutes. It takes 440 minutes for 1 person. It takes 220 minutes for 2 people. In general it takes n people will take 440/n minutes in other words, time= 440/n or n*time=440

OpenStudy (anonymous):

Thanks!!!